# WeatherDiffusion: Controllable Weather Editing in Intrinsic Space

### v2

@@ 1 Introduction
For autonomous vehicles, robust scene understanding requires the ability to handle adverse weather conditions. Techniques that handle weather effects, can substantially improve the robustness of perception models across different weather conditions. While diffusion-based image editing techniques [labs2025flux1kontextflowmatching, wu2025qwenimagetechnicalreport, brooks2023instructpix2pix, hertz2022prompttopromptimageeditingcross] have created significant opportunities here, a key limitation remains: the lack of fine-grained controllability in the generated scenarios.
@@
Following existing image editing approaches, weather editing methods [cosne2020using, li2021weatherganmultidomainweather, tasnim2025normalizing] typically leverage a unified generative model to perform weather transformations in the pixel space. Particularly, WeatherWeaver [lin2025controllable] decomposes weather editing into a two-stage strategy: weather removal and weather synthesis. While these methods have achieved impressive results in weather manipulation, they still struggle to preserve the underlying material and geometry of the scene while generating natural illumination.
@@
In this paper, we present WeatherDiffusion, a novel framework for controllable weather editing that operates in intrinsic space. Our key insight is that by decomposing a scene into its fundamental components—material properties, geometry, and lighting, we can achieve higher control over weather effects than in the image space. WeatherDiffusion consists of two key components: a forward renderer that uses CLIP-space interpolation and diffusion priors for fine-grained weather synthesis, and a novel inverse renderer tailored for outdoor and autonomous driving scenes. We are inspired by recent diffusion-based intrinsic decomposition and recomposition approaches, RGBX [zeng2024rgb] and DiffusionRenderer [liang2025diffusionrenderer], which have achieved impressive results at the indoor and object levels; however, they do not focus on weather and do not generalize to large-scale autonomous driving scenarios. We bridge this gap with a novel intrinsic map-aware attention mechanism that ensures spatial correspondence, thereby significantly enhancing decomposition fidelity in complex, unconstrained outdoor environments. By combining the inverse and forward renderers, our framework allows for controlled editing of weather and lighting.
@@
Our WeatherDiffusion outperforms existing state-of-the-art inverse and forward rendering methods, achieving over 10 dB PSNR improvement in inverse rendering and higher PickScore [kirstain2023pickapicopendatasetuser] in forward rendering. Compared with weather restoration and pixel-space editing methods, our approach achieves higher text-image consistency and better DINO-based structural alignment, enabling cleaner and more controllable weather editing. By proactively correcting environmental distortions at the visual input level, we significantly boost the performance of downstream tasks such as object detection and semantic segmentation. We observe that after applying WeatherDiffusion, the detection and segmentation performance on the ACDC [sakaridis2021acdc] benchmark increases by 87.1% ( from 13.15% to 24.60%) and 24.5% ( from 24.13% to 30.05%), respectively. We also introduce a synthetic dataset and a real-world dataset containing 38k and 18k images with intrinsic map annotations, covering diverse weather conditions and a wide range of driving scenes to train these two components. To summarize, our contributions are as follows:
@@ 2 Related work
Diffusion models. Diffusion models have achieved remarkable progress in high-fidelity and text-conditioned image generation [ho2020denoising, rombach2022high, ramesh2022hierarchical, esser2024scaling]. Modern diffusion frameworks typically adopt a denoising process [ho2020denoising, song2020denoising, liu2022pseudo, lipman2022flow], parameterized by UNet- or DiT-based backbones [ronneberger2015unet, peebles2023dit]. In our research, we repurpose diffusion models to jointly estimate material, geometry, and lighting from an image while synthesizing new images under specified weather conditions. This demonstrates that the strong priors embedded in pre-trained diffusion models can be effectively leveraged for faithful and physically grounded estimations.
@@
The emergence of the diffusion model has catalyzed a novel methodology that leverages generative models to learn the joint probability distribution between images and their corresponding intrinsic maps [kocsis2024intrinsic, zeng2024rgb, li2024idarb, fu2024geowizard, chen2024uni, liang2025diffusionrenderer, he2025unirelightlearningjointdecomposition, IntrinsicDiffusion]. IID [kocsis2024intrinsic] focuses on material estimation in indoor scenes, while RGBX [zeng2024rgb] extends diffusion to bidirectional mapping between RGB images and maps. Several works further incorporate geometric priors [fu2024geowizard] or multi-view cues [li2024idarb]. DiffusionRenderer [liang2025diffusionrenderer] adapts a video diffusion model to achieve temporally consistent inverse and forward rendering, and UniRelight [he2025unirelightlearningjointdecomposition] jointly estimates albedo and relighted video frames. These methods are primarily designed for indoor scenes, small objects, or video relighting, and they struggle to generalize to large-scale outdoor driving scenes with diverse weather conditions. Moreover, none of these works address controllable weather editing or decomposition in intrinsic space under multiple weather conditions, which is the goal of our framework.
@@
Most image editing techniques operate purely in the pixel space, using diffusion models to modify color, texture, or local appearance without modeling underlying scene factors [tumanyan2022plugandplaydiffusionfeaturestextdriven, brooks2023instructpix2pix, wu2025qwenimagetechnicalreport, labs2025flux1kontextflowmatching, hertz2022prompttopromptimageeditingcross, cao2023masactrltuningfreemutualselfattention, parmar2023zeroshotimagetoimagetranslation, kawar2023imagictextbasedrealimage]. Following this paradigm, existing weather-editing models [cosne2020using, li2021weatherganmultidomainweather, tasnim2025normalizing] directly translate one weather type to another, and WeatherWeaver [lin2025controllable] finetunes a video diffusion model for weather removal and synthesis. However, pixel-space editing lacks physical interpretability and cannot guarantee consistent material, geometry, or illumination. Intrinsic-space manipulation has been explored by IntrinsicEdit [Lyu_2025], but it focuses on object-level editing rather than large-scale outdoor scenes. Other approaches edit weather in 3D space [qian2025weathereditcontrollableweatherediting, li2023climatenerfextremeweathersynthesis], but they require accurate geometry, which is unavailable for real-world driving scenes. In contrast, WeatherDiffusion performs controllable weather editing in the intrinsic space. This formulation enables fine-grained and geometry-preserving control that is difficult to achieve in the pixel space.
@@ 3.1 WeatherDiffusion
We repurpose Stable Diffusion 3.5 [sd3.5_medium] to enable the inverse and forward renderers. To leverage the diffusion prior to achieve fine-grained weather control, we first obtain a weather transitional direction in the CLIP space:
@@ 3.2 Intrinsic map-aware attention
We identify important regions based on the intrinsic maps and use them to construct an attention bias. Specifically, we first employ DINOv2 [oquab2023dinov2] to extract a set of patch tokens . For each intrinsic map, we define a learnable embedding that captures its inherent characteristics. A gating mechanism is applied to selectively filter patch tokens based on the current intrinsic map. Formally, we compute a map-aware mask:
@@ 3.3 Dataset construction
Existing datasets of images with corresponding intrinsic maps suffer from the absence of outdoor environments (OpenRooms [li2020inverse], Hypersim [roberts2021hypersim], InteriorVerse [zhu2022learning]) or insufficient weather diversity (MatrixCity [li2023matrixcity]), and thus are inappropriate for large-scale outdoor and autonomous driving scenes. While MatrixCity provides continuous variations of fog density and illumination, it mainly targets relighting and reconstruction tasks. In contrast, our work requires diverse weather and lighting conditions rather than smooth transitions of the same type. Moreover, MatrixCity suffers from legal licensing issues. To fill the gap in large-scale weather-diverse autonomous driving datasets with paired images and intrinsic maps, we propose WeatherSynthetic and WeatherReal. Sample images from our datasets are shown in Fig. 5.
@@
is a real-world dataset on autonomous driving scenes with various weather conditions. We use our inverse renderer to generate intrinsic maps of open-source datasets such as Waymo [sun2020scalability] and Kitti [geiger2013vision]. We use a multimodal model to remove challenging scenarios (e.g., rainy nights) to ensure the generation of high-quality pseudo-labels, and check the quality manually. The samples are shown in the supplementary materials. Moreover, we employ a pre-trained image editing model (i.e., Flux-Kontext [labs2025flux1kontextflowmatching]) to alter the weather types. Our WeatherReal is motivated by the observation that after training our model merely on synthetic data, the inverse and forward renderers lack sufficient generalization capability on real-world samples. Note that WeatherReal is only used to finetune models and is not used to evaluate results.
@@ 4 Experimental results
Following WeatherWeaver [lin2025controllable], we use PickScore [kirstain2023pickapicopendatasetuser], CLIP image-text consistency (denoted as CLIP-S), and DINO structure similarity (denoted as DINO-S) to evaluate editing results. Following previous works [zeng2024rgb, li2024idarb], we report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Mean Angular Error (MAE), and Learned Perceptual Image Patch Similarity (LPIPS) for inverse rendering. We compare our performance with pixel-space editing methods (Flux-Kontext [labs2025flux1kontextflowmatching], Qwen-Image-Edit [wu2025qwenimagetechnicalreport], Instruct-Pix2Pix [brooks2023instructpix2pix], WeatherWeaver [lin2025controllable]) and weather restoration methods (AWRaCL [rajagopalan2025awracle], Histoformer [sun2024restoringimagesadverseweather]). We also compare our inverse and forward rendering results with RGBX [zeng2024rgb], IID [kocsis2024intrinsic], Geowizard [fu2024geowizard], IDArb [li2024idarb] and DiffusionRenderer [liang2025diffusionrenderer]. We evaluate different methods on WeatherSynthetic, Waymo [sun2020scalability], TransWeather [valanarasu2021transweather], ACDC [sakaridis2021acdc], and additional Internet images covering diverse weather conditions.
@@ 4.1 Comparison with pixel-space editing methods
We show quantitative comparisons in Tab. 1. Our method achieves the highest CLIP-S, indicating that it produces the most text-aligned and plausible weather editing results. In terms of DINO-S, we rank second only to Flux-Kontext [labs2025flux1kontextflowmatching], which, however, fails to remove or synthesize weather effects effectively. Although Qwen-Image-Edit [wu2025qwenimagetechnicalreport] achieves a slightly higher PickScore, it often introduces inconsistent textures and geometry. PickScore is suitable for measuring user preference, but it does not measure physical consistency and editing plausibility.
@@
| Method | PickScore [kirstain2023pickapicopendatasetuser] | CLIP-S | DINO-S | |||
@@ 4.2 Comparison with weather restoration methods
We present qualitative comparisons with AWRaCLe [rajagopalan2025awracle] and Histoformer [sun2024restoringimagesadverseweather] in Fig. 8. The advantage of our method is its ability to effectively remove various degradations in adverse weather images, such as airborne particles (like snowflakes), snow on the ground, and overall illumination. The weather restoration methods only remove particles while failing to change surface material or lighting conditions.
To further verify the improvement brought by different models to downstream applications, we choose object detection and semantic segmentation as tasks, evaluating performance improvement before and after weather editing. Specifically, we chose DETR [carion2020detr] and Segformer [xie2021segformer] as base models. When applying object detection and semantic segmentation models to our re-rendered images, both tasks achieve more accurate and consistent results, as shown in the lower rows of Fig. 8. Following this way, we apply weather editing and evaluate on the validation set of the ACDC benchmark [sakaridis2021acdc], and the results are shown in Tab. 2.
@@ 4.3.1 Inverse rendering
In this part, we first conduct quantitative evaluations on WeatherSynthetic, then we evaluate on real-world TransWeather [valanarasu2021transweather] datasets. We show the comparison between our method and existing methods on the test set of our WeatherSynthetic in Tab. 3. Our method outperforms existing approaches across all evaluation metrics. We fine-tune IID [kocsis2024intrinsic] and RGBX [zeng2024rgb] with the same training steps on our WeatherSynthetic. Their performance improves, but they fail to provide high-quality estimation. An overall qualitative comparison is shown in the supplementary.
@@ 4.3.2 Forward rendering
A comparison of forward rendering between WeatherDiffusion, RGBX, and DiffusionRenderer [liang2025diffusionrenderer] is shown in Fig. 10. We use the prompt “A sunny day in the city.” for ours and RGBX, and provide an environment lighting for DiffusionRenderer. Our WeatherDiffusion produces images that better align with the text description than RGBX, and avoid abnormal textures and illuminations. DiffusionRenderer fails to recover all details, such as road signs and distant buildings. As shown in Tab. 1, our method achieves the highest metrics in rendering-based methods on all weather conditions. More results are shown in the supplementary.
@@ 4.5 Limitations
Our framework is designed for single-image weather editing in the intrinsic space. Temporal modeling introduces additional factors, such as object motion and occlusion changes, that are orthogonal to our core contribution and require a video prior. Therefore, the current framework does not guarantee temporal consistency in video sequences, as shown in Fig. 12. Extending WeatherDiffusion to videos would likely be possible by building on a video diffusion model, similar to DiffusionRenderer [liang2025diffusionrenderer]. However, these models typically demand substantially more training data, computational resources, and operate at lower resolutions, making them difficult to apply to driving scenes directly. Building a temporally consistent framework is an important direction for future work, but it remains beyond the scope of this paper.

### v3

@@ 1 Introduction
For autonomous vehicles, robust scene understanding requires the ability to handle adverse weather conditions. Techniques that handle weather effects, can substantially improve the robustness of perception models across different weather conditions. While diffusion-based image editing techniques [16, 44, 2, 11] have created significant opportunities here, a key limitation remains: the lack of fine-grained controllability in the generated scenarios.
@@
Following existing image editing approaches, weather editing methods [6, 17, 41] typically leverage a unified generative model to perform weather transformations in the pixel space. Particularly, WeatherWeaver [23] decomposes weather editing into a two-stage strategy: weather removal and weather synthesis. While these methods have achieved impressive results in weather manipulation, they still struggle to preserve the underlying material and geometry of the scene while generating natural illumination.
@@
In this paper, we present WeatherDiffusion, a novel framework for controllable weather editing that operates in intrinsic space. Our key insight is that by decomposing a scene into its fundamental components—material properties, geometry, and lighting, we can achieve higher control over weather effects than in the image space. WeatherDiffusion consists of two key components: a forward renderer that uses CLIP-space interpolation and diffusion priors for fine-grained weather synthesis, and a novel inverse renderer tailored for outdoor and autonomous driving scenes. We are inspired by recent diffusion-based intrinsic decomposition and recomposition approaches, RGBX [46] and DiffusionRenderer [22], which have achieved impressive results at the indoor and object levels; however, they do not focus on weather and do not generalize to large-scale autonomous driving scenarios. We bridge this gap with a novel intrinsic map-aware attention mechanism that ensures spatial correspondence, thereby significantly enhancing decomposition fidelity in complex, unconstrained outdoor environments. By combining the inverse and forward renderers, our framework allows for controlled editing of weather and lighting.
@@
Our WeatherDiffusion outperforms existing state-of-the-art inverse and forward rendering methods, achieving over 10 dB PSNR improvement in inverse rendering and higher PickScore [14] in forward rendering. Compared with weather restoration and pixel-space editing methods, our approach achieves higher text-image consistency and better DINO-based structural alignment, enabling cleaner and more controllable weather editing. By proactively correcting environmental distortions at the visual input level, we significantly boost the performance of downstream tasks such as object detection and semantic segmentation. We observe that after applying WeatherDiffusion, the detection and segmentation performance on the ACDC [37] benchmark increases by 87.1% ( from 13.15% to 24.60%) and 24.5% ( from 24.13% to 30.05%), respectively. We also introduce a synthetic dataset and a real-world dataset containing 38k and 18k images with intrinsic map annotations, covering diverse weather conditions and a wide range of driving scenes to train these two components. To summarize, our contributions are as follows:
@@ 2 Related work
Diffusion models. Diffusion models have achieved remarkable progress in high-fidelity and text-conditioned image generation [12, 35, 33, 7]. Modern diffusion frameworks typically adopt a denoising process [12, 38, 25, 24], parameterized by UNet- or DiT-based backbones [36, 30]. In our research, we repurpose diffusion models to jointly estimate material, geometry, and lighting from an image while synthesizing new images under specified weather conditions. This demonstrates that the strong priors embedded in pre-trained diffusion models can be effectively leveraged for faithful and physically grounded estimations.
@@
The emergence of the diffusion model has catalyzed a novel methodology that leverages generative models to learn the joint probability distribution between images and their corresponding intrinsic maps [15, 46, 21, 8, 5, 22, 10, 26]. IID [15] focuses on material estimation in indoor scenes, while RGBX [46] extends diffusion to bidirectional mapping between RGB images and maps. Several works further incorporate geometric priors [8] or multi-view cues [21]. DiffusionRenderer [22] adapts a video diffusion model to achieve temporally consistent inverse and forward rendering, and UniRelight [10] jointly estimates albedo and relighted video frames. These methods are primarily designed for indoor scenes, small objects, or video relighting, and they struggle to generalize to large-scale outdoor driving scenes with diverse weather conditions. Moreover, none of these works address controllable weather editing or decomposition in intrinsic space under multiple weather conditions, which is the goal of our framework.
@@
Most image editing techniques operate purely in the pixel space, using diffusion models to modify color, texture, or local appearance without modeling underlying scene factors [42, 2, 44, 16, 11, 3, 29, 13]. Following this paradigm, existing weather-editing models [6, 17, 41] directly translate one weather type to another, and WeatherWeaver [23] finetunes a video diffusion model for weather removal and synthesis. However, pixel-space editing lacks physical interpretability and cannot guarantee consistent material, geometry, or illumination. Intrinsic-space manipulation has been explored by IntrinsicEdit [27], but it focuses on object-level editing rather than large-scale outdoor scenes. Other approaches edit weather in 3D space [31, 19], but they require accurate geometry, which is unavailable for real-world driving scenes. In contrast, WeatherDiffusion performs controllable weather editing in the intrinsic space. This formulation enables fine-grained and geometry-preserving control that is difficult to achieve in the pixel space.
@@ 3.1 WeatherDiffusion
We repurpose Stable Diffusion 3.5 [1] to enable the inverse and forward renderers. To leverage the diffusion prior to achieve fine-grained weather control, we first obtain a weather transitional direction in the CLIP space:
@@ 3.2 Intrinsic map-aware attention
We identify important regions based on the intrinsic maps and use them to construct an attention bias. Specifically, we first employ DINOv2 [28] to extract a set of patch tokens . For each intrinsic map, we define a learnable embedding that captures its inherent characteristics. A gating mechanism is applied to selectively filter patch tokens based on the current intrinsic map. Formally, we compute a map-aware mask:
@@ 3.3 Dataset construction
Existing datasets of images with corresponding intrinsic maps suffer from the absence of outdoor environments (OpenRooms [20], Hypersim [34], InteriorVerse [47]) or insufficient weather diversity (MatrixCity [18]), and thus are inappropriate for large-scale outdoor and autonomous driving scenes. While MatrixCity provides continuous variations of fog density and illumination, it mainly targets relighting and reconstruction tasks. In contrast, our work requires diverse weather and lighting conditions rather than smooth transitions of the same type. Moreover, MatrixCity suffers from legal licensing issues. To fill the gap in large-scale weather-diverse autonomous driving datasets with paired images and intrinsic maps, we propose WeatherSynthetic and WeatherReal. Sample images from our datasets are shown in Fig. 5.
@@
is a real-world dataset on autonomous driving scenes with various weather conditions. We use our inverse renderer to generate intrinsic maps of open-source datasets such as Waymo [39] and Kitti [9]. We use a multimodal model to remove challenging scenarios (e.g., rainy nights) to ensure the generation of high-quality pseudo-labels, and check the quality manually. The samples are shown in the supplementary materials. Moreover, we employ a pre-trained image editing model (i.e., Flux-Kontext [16]) to alter the weather types. Our WeatherReal is motivated by the observation that after training our model merely on synthetic data, the inverse and forward renderers lack sufficient generalization capability on real-world samples. Note that WeatherReal is only used to finetune models and is not used to evaluate results.
@@ 4 Experimental results
Following WeatherWeaver [23], we use PickScore [14], CLIP image-text consistency (denoted as CLIP-S), and DINO structure similarity (denoted as DINO-S) to evaluate editing results. Following previous works [46, 21], we report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Mean Angular Error (MAE), and Learned Perceptual Image Patch Similarity (LPIPS) for inverse rendering. We compare our performance with pixel-space editing methods (Flux-Kontext [16], Qwen-Image-Edit [44], Instruct-Pix2Pix [2], WeatherWeaver [23]) and weather restoration methods (AWRaCL [32], Histoformer [40]). We also compare our inverse and forward rendering results with RGBX [46], IID [15], Geowizard [8], IDArb [21] and DiffusionRenderer [22]. We evaluate different methods on WeatherSynthetic, Waymo [39], TransWeather [43], ACDC [37], and additional Internet images covering diverse weather conditions.
@@ 4.1 Comparison with pixel-space editing methods
We show quantitative comparisons in Tab. 1. Our method achieves the highest CLIP-S, indicating that it produces the most text-aligned and plausible weather editing results. In terms of DINO-S, we rank second only to Flux-Kontext [16], which, however, fails to remove or synthesize weather effects effectively. Although Qwen-Image-Edit [44] achieves a slightly higher PickScore, it often introduces inconsistent textures and geometry. PickScore is suitable for measuring user preference, but it does not measure physical consistency and editing plausibility.
@@
| Method | PickScore [14] | CLIP-S | DINO-S | |||
@@ 4.2 Comparison with weather restoration methods
We present qualitative comparisons with AWRaCLe [32] and Histoformer [40] in Fig. 8. The advantage of our method is its ability to effectively remove various degradations in adverse weather images, such as airborne particles (like snowflakes), snow on the ground, and overall illumination. The weather restoration methods only remove particles while failing to change surface material or lighting conditions.
To further verify the improvement brought by different models to downstream applications, we choose object detection and semantic segmentation as tasks, evaluating performance improvement before and after weather editing. Specifically, we chose DETR [4] and Segformer [45] as base models. When applying object detection and semantic segmentation models to our re-rendered images, both tasks achieve more accurate and consistent results, as shown in the lower rows of Fig. 8. Following this way, we apply weather editing and evaluate on the validation set of the ACDC benchmark [37], and the results are shown in Tab. 2.
@@ 4.3.1 Inverse rendering
In this part, we first conduct quantitative evaluations on WeatherSynthetic, then we evaluate on real-world TransWeather [43] datasets. We show the comparison between our method and existing methods on the test set of our WeatherSynthetic in Tab. 3. Our method outperforms existing approaches across all evaluation metrics. We fine-tune IID [15] and RGBX [46] with the same training steps on our WeatherSynthetic. Their performance improves, but they fail to provide high-quality estimation. An overall qualitative comparison is shown in the supplementary.
@@ 4.3.2 Forward rendering
A comparison of forward rendering between WeatherDiffusion, RGBX, and DiffusionRenderer [22] is shown in Fig. 10. We use the prompt “A sunny day in the city.” for ours and RGBX, and provide an environment lighting for DiffusionRenderer. Our WeatherDiffusion produces images that better align with the text description than RGBX, and avoid abnormal textures and illuminations. DiffusionRenderer fails to recover all details, such as road signs and distant buildings. As shown in Tab. 1, our method achieves the highest metrics in rendering-based methods on all weather conditions. More results are shown in the supplementary.
@@ 4.5 Limitations
Our framework is designed for single-image weather editing in the intrinsic space. Temporal modeling introduces additional factors, such as object motion and occlusion changes, that are orthogonal to our core contribution and require a video prior. Therefore, the current framework does not guarantee temporal consistency in video sequences, as shown in Fig. 12. Extending WeatherDiffusion to videos would likely be possible by building on a video diffusion model, similar to DiffusionRenderer [22]. However, these models typically demand substantially more training data, computational resources, and operate at lower resolutions, making them difficult to apply to driving scenes directly. Building a temporally consistent framework is an important direction for future work, but it remains beyond the scope of this paper.